At 125 degrees C, the equilibrium constant Kp for the following reaction is 0.25.

2 NaHCO3 (s) \leftrightharpoons Na2CO3(s) + CO2(g) + H2O(g)

a.) Calculate Kc for the reaction at 125 degrees C

b.) At the beginning of the reaction given in part (A) only 10.0g of NaHCO3 was there in an evacuated 1.00L flask. What is the total pressure in the flask, after the equilibrium is established 125 degrees C?
Answer #1

Ans The equillibiwn equation is written as. Nahco(s) = Nagco,CS) + COgGg) + H30 Cg) ke for the reaction is 0.25 Relation betwPartial fuegs wie at equillo an (Here, Pr = Total pressure at eqwillibaium. ) Kp = (2) (DPP 0.25 = 5 (P2) (P, 22- 1.00 1 Pp =